the length of a rectangle is 3 in more than twice the width. the area is 189^2 what is the length of the triangle?
Answer by mananth(16946)

the length of a rectangle is 3 in more than twice the width. the area is 189 what is the length of the rectangle?
let the width be x
length = 2x+3
x(2x+3) = 189
2x^2+3x-189=0
2x^2+-18x+21x-189=0
2x(x-9)+21(x-9)=0
(x-9)(2x+21)=0
x=9 or x= -21/2
x=9 which is the width since the width can't be negative